difficulty with these two back to back questions. 41-42.

Suppose the security i has the following probability distributions: Stato Probability 0.3 0.7 -10% 24% What is the expected return for security i? 5.6% 8.296 90% o 10.4% 13.8% QUESTION 42 What is the standard deviation for security i? 9.84% О 10.58% 13.08% 15.58% Click Save and Submit to save and submit. Click Save All Answers to save all answers acBook Pro
Answer #1

Expected return = P1*R1 + P2*R2 = 0.3*-10% + 0.7*24%

= 13.8%

Variance= P1*(R1-ER)^2+ P2*(R1-ER)^2

=0.3* (-10%-13.8%)^2 + 0.7*(24%-13.8%)^2

=0.024276

Standard deviation= sqrt* Variance

= SQRT (0.024276)

=15.58%